Distance between two parallel planes 2x + y + 2z = 8 and 4x + 2y + 4z + 5 = 0 is
(1)
5/2
(2)
7/2
(3)
9/2
(4)
3/2

Grade:upto college level

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Given planes are2x + y + 2z = 8 and 4x + 2y + 4z + 5 = 0
multiplying 1stequation by 2 we get
4x+2y+4z=16 and 4x+2y+4z=-5
distance between the planes=16-(-5)/sqrt(42+22+42)
=21/6=7/2
Thanks and Regards
Shaik Aasif
askIITians faculty
ankit singh
3 years ago
|c1-c2|\\sqrt{a2=b2=c2}simple firstly make x,y,z cofficient of both planes equation in same ratio thansolve easily thank u regard ankit singh